A solution containing 8.75 ppm KMnO4 has transmittance of 0.743 in a 1.00 cm cell at 520 nm.
Answer #1

concentration(C) = 8.75 ppm = 0.00875 gm /L = 5.53*10-5 mole /L = 5.53*10-5 M

[ molar mass of KMnO4 = 158.034 gm and 0.00875 gm = 5.53*10-5 mole]

transmittance = 0.743

so, absorbance (A) = - log T = - log(0.743) = 0.1290

path length (l) = 1 cm

according to Lamber Beers law

A =\varepsilon cl [ where \varepsilon molar absorptivity]

or, \varepsilon = 2332.7 M-1 cm-1